44. Geometry. The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by malakumar_kos@yahoo.com(315)


. The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field
Let the width of the rectanglular field be = x ft
Then length of the rectanglular field = 2x-5 ft
perimeter of the field is = 230 ft
perimeter = 2(length+width)
230 = 2(2x-5+x)
230/2 = 3x-5
115 +5 = 3x
120 = 3x
120/3 = x
x = 40 ft
width of the rectangular field = 40 ft
Length of the rectangular field = 75 ft (2*40-5= 80-5=75)